# Collapsing in polygonal dynamics

### v2

@@ Abstract
Polygonal dynamics is a family of dynamical systems containing many studied systems, like the famous pentagram map. Similar collapsing phenomena seem to occur in most of these systems. We give a unifying, general definition of polygonal dynamics, and conjecture that a generic orbit collapses towards a predictable point. We manage to prove it in some setting. For the special case of “closed polygons”, we show as a corollary that the collapse point depends algebraicly on the vertices of the starting polygon, using tools called scaling symmetry and infinitesimal monodromy. This generalizes previous results about the pentagram map. Then, we investigate the case of polygonal dynamics in $\mathbb{P}^{1}$ for which we give an explicit polynomial equation satisfied by the collapse point. Based on previous works, we define a new dynamical system, the “staircase” cross-ratio dynamics, for which we study particular configurations.
@@ 1 Introduction
Let $n,\penalty 10000\ d$ be positive integers. A twisted $\bm{n}$-gon $P$ is the data of $(p_{1},\dots,p_{n})$ and $M$, where the $p_{i}\in\mathbb{P}^{d}$ are the vertices and $M\in\mathbb{P}\mathrm{GL}_{d+1}$ is the monodromy. The vertices are to be thought as an infinite sequence satisfying the condition:
@@
$P$ is said to be closed if $M=\mathrm{Id}$. The set of twisted and closed $n$-gons are respectively denoted by $\tilde{\mathcal{T}}_{n}$ and $\tilde{\mathcal{P}}_{n}$.
A polygonal transformation is a birational map $T:\tilde{\mathcal{T}}_{n}\dashrightarrow\tilde{\mathcal{T}}_{n}$ which preserves the monodromy and commutes with the following action of $\mathbb{P}\mathrm{GL}_{d+1}$:
@@
The class of $P$ is denoted by $[P]$. The action of $\mathbb{P}\mathrm{GL}_{d+1}$ allows to define the quotient spaces [^2] Sometimes taking a quotient is more tedious, and we need to use geometric invariant theory (GIT). See [1, §3.1] for details. $\mathcal{T}_{n}$ and $\mathcal{P}_{n}$, also called moduli spaces.
Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$, it also yields a dynamic over the quotient spaces. As we will see later, it is often very different from the one on the initial space.
@@
We denote the projective spaces by $\mathbb{P}^{d}$ and their automorphism groups by $\mathbb{P}\mathrm{GL}_{d+1}$ without specifying over which field $k$ they are defined. This is because we see them from the algebraic geometry standpoint, as schemes (like it’s done in [1]). Similarly, the group scheme $\mathbb{G}_{m}$ refers to the group of invertible elements $k^{*}$. However, no algebraic geometry background will be needed here, and the unfamiliar reader may consider the statements over a fixed field, like $\mathbb{C}$.
@@
This definition can be generalized by replacing $\mathbb{P}^{d}$ by a topological space $X$ and $\mathbb{P}\mathrm{GL}_{d+1}$ by a group $G$ acting by homeomorphism (see [2, §2]). For instance, the polygon recutting defined by Adler [3] in 1993 takes place in $X=\mathbb{R}^{2}$ with $G=\mathrm{Isom}^{+}(\mathbb{R}^{2})$. See table LABEL:tab:examples_polygonal_dynamics for a review of existing polygonal dynamics.
Since our results only concerns polygonal dynamics in $\mathbb{P}^{d}$, we will set the rest aside for now.
Figure 1: The convex pentagon $P$, lying on the real plane, is sent to $P^{\prime}$ by the pentagram map.
Let’s start with the most famous polygonal dynamic: the pentagram map, defined by Schwartz [4] in 1992. In his original paper, he defines the map $T$ acting on the set of strictly convex closed $n$-gons (with $n\geq 5$, hence the name) in $\mathbb{P}^{2}(\mathbb{R})$. [^3] He generalised it to twisted polygons [5] in 2007. A polygon $P=(p_{1},\dots,p_{n})$ is sent to another one by taking the intersections of the small diagonals $\overline{p_{i-1}p_{i+1}}$ and $\overline{p_{i}p_{i+2}}$, where the indices are understood modulo $n$ (see figure 1 for an example). Schwartz proved that for these real and strictly convex polygons, the sequence of iterates collapses exponentially fast towards a point. However, the meaning of this point, or even the existence of a formula to express it from the original vertices of $P$, remained unknown for 26 years. This collapse point, which can be thought as a kind of “projective center of mass” of the polygon $P$, is quite intriguing.
In 2018, Glick [6] gave a formula for the collapse point. To do this, he defines a linear operator associated to a convex polygon $P$, now called Glick’s operator, which is invariant by the dynamic and sends $P$ to its interior. He concludes that the collapse point must be a fixed point of the operator. The formula is quite simple, involving only a cubic extension of the field generated over $\mathbb{Q}$ by the coordinates of the vertices of $P$. However, the proof is quite computational and doesn’t shed much light on the meaning of the point.
In 2020, Aboud and Izosimov [7] reinterpret this point as being an eigenline of a certain $3\times 3$ matrix $M^{\prime}$ associated to the polygon $P$, called its infinitesimal monodromy. This matrix can be in some way identified with Glick’s operator. To do this, they “open” the closed polygon into a family of twisted ones through a tool called the scaling symmetry.

### v3

@@ Abstract
We define polygonal dynamics as a family of dynamical systems acting on points in projective spaces. The most famous example is the pentagram map. Similar collapsing phenomena seem to occur in most of these systems. We prove it in some case, and conjecture that it almost always happens. Moreover, we give a formula for the limit point in term of roots of $d+1$ degree polynomials (where $d$ is the dimension of the projective space). We do so by generalizing Glick’s operator, interpreted as an infinitesimal monodromy. This answers questions about its reappearance in many systems, together with preserved quantities. We apply these results to several polygonal dynamics in $\mathbb{P}^{1}$ and introduce a new one called “staircase” cross-ratio dynamics, for which we study particular configurations.
@@ 1 Introduction
Before starting, a word on vocabulary. We denote the projective spaces by $\mathbb{P}^{d}$ and their automorphism groups by $\mathbb{P}\mathrm{GL}_{d+1}$ without specifying over which field $F$ they are defined. This is because we see them from the algebraic geometry standpoint, as schemes (like it’s done in [weinreichAlgebraicDynamicsPentagram2023]). Similarly, the group scheme $\mathbb{G}_{m}$ refers to the group of invertible elements $F^{*}$. However, no algebraic geometry background will be needed here, and the unfamiliar reader may consider the statements over a fixed field, like $\mathbb{C}$.
1.1 Definitions
@@
Let $n,d$ be positive integers. A twisted $\bm{n}$-gon $P$ is the data of vertices $p_{1},\dots,p_{n}\in\mathbb{P}^{d}$ and a monodromy $M\in\mathbb{P}\mathrm{GL}_{d+1}$. The vertices are to be thought as an infinite sequence satisfying:
@@
Moreover, we impose a non-degeneracy condition, meaning that for any $i\in\mathbb{Z}$, the vertices $(p_{i},\dots,p_{i+d+1})$ form a projective frame.
$P$ is said be closed if $M=\mathrm{Id}$. The set of twisted and closed $n$-gons are respectively denoted by $\tilde{\mathcal{T}}_{n}$ and $\tilde{\mathcal{P}}_{n}$.
A polygonal transformation is a birational map
$T:\tilde{\mathcal{T}}_{n}\dashrightarrow\tilde{\mathcal{T}}_{n}$
which commutes with the following action of $\mathbb{P}\mathrm{GL}_{d+1}$:
@@
The class of $P$ is denoted by $[P]$. The action of $\mathbb{P}\mathrm{GL}_{d+1}$ allows to define the quotient spaces $\mathcal{T}_{n}$ and $\mathcal{P}_{n}$, also called moduli spaces.
Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$, it also yields a birational map (again denoted by $T$) over the moduli spaces. In many cases, the dynamic of $T$ on the moduli spaces have a very strong property called discrete integrability. [^2] The following definition works for the algebraic case, and is inspired from [weinreichAlgebraicDynamicsPentagram2023]. For general review of continuous integrable systems, and the link between the Hamiltonian and algebraic flavors, see [audin:hal-00135277].
Definition 1.2.
A dynamical system $(X,T)$ is said to be discretely integrable if there is a family $A$ of abelian varieties [^3] Here, this is the algebraic counterpart of lagrangian tori. which foliate $X$ such that, on a dense subset $\Sigma\subset X$, there is a birational map $\delta:\Sigma\to A$ which identifies some iterate of $T$ with a relative translation $\tau:A\to A$.
@@
Definition 1.1 can be generalized by replacing $\mathbb{P}^{d}$ with a topological space $X$ and $\mathbb{P}\mathrm{GL}_{d+1}$ by a group $G$ acting by homeomorphism (see [beffaIntegrableGeneralizationsPentagram2015, §2]). For instance, the polygon recutting defined by Adler [adlerRecuttingsPolygons1993] in 1993 takes place in $X=\mathbb{R}^{2}$ with $G=\mathrm{Isom}^{+}(\mathbb{R}^{2})$. See table LABEL:tab:examples_polygonal_dynamics for a review of existing polygonal dynamics. Since our results only concerns polygonal dynamics in $\mathbb{P}^{d}$, we will set the rest aside.
1.2 Previous works
Figure 1: The pentagram map applied on a closed pentagon in $\mathbb{P}^{2}(\mathbb{R})$.
Let’s start with the most famous polygonal dynamic: the pentagram map, defined by Schwartz [schwartzPentagramMap1992] in 1992. In his original paper, he defines the map $T$ acting on the set of strictly convex closed $n$-gons (with $n\geq 5$, hence the name) in $\mathbb{P}^{2}(\mathbb{R})$. [^4] He generalised it to twisted polygons [schwartzDiscreteMonodromyPentagrams2007] in 2007. A polygon $P=(p_{1},\dots,p_{n})$ is sent to another one by taking the intersections of the small diagonals $\overline{p_{i-1}p_{i+1}}$ and $\overline{p_{i}p_{i+2}}$, where the indices are understood modulo $n$ (see figure 1 for an example). Schwartz proved that for strictly convex polygons, the sequence of iterates collapses exponentially fast towards a point. However, the meaning of this point, or even the existence of a formula to express it from the original vertices of $P$, remained unknown for 26 years. This collapse point, which can be thought as a kind of “projective center of mass” of the polygon $P$, is quite intriguing.
In 2018, Glick [glickLimitPointPentagram2020] gives a formula for the collapse point. To do this, he defines a linear operator of $\mathbb{R}^{3}$ associated to a convex polygon $P$, now called Glick’s operator, which is invariant by the dynamic and sends $P$ to its interior. He concludes that the collapse point must be a fixed point of the operator. The formula is quite simple, involving polynomials of degree $3$ whose coefficients depend on the vertices of $P$. However, the proof is quite computational and doesn’t shed much light on the meaning of the point.
In 2020, Aboud and Izosimov [aboudLimitPointPentagram2022] reinterpret Glick’s operator in term of a matrix $M^{\prime}$, called an infinitesimal monodromy. To do this, they “open” the closed polygon into a family of twisted ones through a tool called a scaling symmetry.
Definition 1.3.
A scaling symmetry is a non-trivial group action
$\rho:\mathbb{G}_{m}\times\mathcal{T}_{n}\to\mathcal{T}_{n}$
which commutes with the dynamic of $T$.
In 2023, Arnold and Arreche [arnoldSymmediansHyperbolicBarycenters2024, §7.2] notice a similarity between the infinitesimal monodromy of the pentagram map and a so-called “hyperbolic barycenter” defined for the flat cross-ratio dynamics (see [arnoldCrossratioDynamicsIdeal2022, §6.2.1]).
In 2024, Schwartz [schwartzFlappingBirdsPentagram2026, §9.2] investigates a family of transformations $\Delta_{k}$ (the pentagram map being $\Delta_{1}$) for which certain polygons collapse. Moreover, he conjectures that the collapsing point is given by a cousin of Glick’s operator. In the first version of his preprint555arXiv:2403.05735v1 page 57., he makes the following observation:
I wonder if this means that the collapse point exists for all starting points of the pentagram map. Even if the iterations go completely crazy under the map, perhaps they still collapse to the point predicted by Glick’s operator. The idea of a completely general collapse point has always seemed absurd to me, but maybe it is not. Nobody knows.
1.3 Plan of the paper and main results
In this paper, we aim to answer these questions by considering the general setting of polygonal dynamics in $\mathbb{P}^{d}$, for any $d\in\mathbb{N}^{*}$. Examples of such systems are presented in Table LABEL:tab:examples_polygonal_dynamics, located in the appendix.
In Section 2, we construct a coordinate system on the moduli space using a generalization of the usual cross-ratio. This allows us to determine the monodromy of any polygon. Then, if the system admits a scaling symmetry, we use it to deform polygons. In particular, we generalize the computation of [aboudLimitPointPentagram2022] and give an explicit formula of the infinitesimal monodromy in term of the cross-ratio coordinates (Theorem 2.10). If the scaling symmetry acts linearly, [^6] Meaning that the cross-ratio coordinates vary linearly, which is the case in every examples. we compute higher derivatives of the monodromy.
In Section 3, we investigate the collapsing of polygons, meaning that all the vertices converge towards the same point. Our computer simulations seem to indicate that it almost always happens when the dynamic takes place over $\mathbb{C}$. We manage to prove it for polygons whose orbit is periodic on the moduli space (Theorem 3.1), which are known to exist. Moreover, we state in Conjecture 3.4 that it is the case for almost every polygons. In Theorem 3.5, we prove that the collapse point of a closed polygon has to be a fixed point of the infinitesimal monodromy, which answers the question of Schwartz in [schwartzFlappingBirdsPentagram2026, §9.2]. We obtain in Corollary 3.6 that the coordinates of the limit point are roots of degree $d+1$ polynomials whose coefficients are determined by the polygon $P$.
In Section 4, we focus on dynamics in $\mathbb{P}^{1}$ and give effective versions of the previous results. In Proposition 4.1, we prove that any dynamic is trivial on the moduli spaces of closed triangles and quadrilaterals. In Theorem 4.2, we give a more explicit formula of the infinitesimal monodromy, which is exactly the matrix which appeared in [arnoldCrossratioDynamicsIdeal2022, §6.2.1]). This answers to the question of Arnold and Arreche [arnoldSymmediansHyperbolicBarycenters2024, §7.2]. We also write explicitly the quadratic polynomial for which the limit point vanishes. We conclude the section by deriving the formulas for two sets of $\left\lfloor n/2\right\rfloor$ preserved quantities (Propositions 4.4 and 4.5), which are good candidates for Hamiltonians.
In Section 5, we apply these results to two polygonal dynamics in $\mathbb{P}^{1}$, namely the leapfrog map (defined in [gekhtmanIntegrableClusterDynamics2016, §5.2]) and the “flat” cross-ratio dynamic (studied in [hetrich-jerominPeriodicDiscreteConformal2001], [arnoldCrossratioDynamicsIdeal2022] and [affolterIntegrableDynamicsProjective2023]). We find their scaling symmetries (Propositions 5.1 and 5.5) and derive their associated infinitesimal monodromy and limit point.
In Section 6, we introduce a new polygonal dynamic on $\mathbb{P}^{1}$, called the “staircase” cross-ratio dynamics. As the name shows, it is a cousin of the “flat” cross-ratio dynamics. They both stem from the theory of discrete holomorphic functions and their continuation (see [affolterSchwarzianOctahedronRecurrence2024, §6] and references therein). They differ by a choice of geometry (flat or staircase like) for the initial conditions. The “staircase” cross-ratio dynamics consist of a group of polygonal transformations called “flips”, which happens to be the affine symmetric group (Theorem 6.4). As for the other dynamics, we find a scaling symmetry and derive its infinitesimal monodromy and limit point. We conclude with a family of $n$-gons with periodic orbits on their moduli space, whose vertices always collapse toward the fixed points of the infinitesimal monodromy, as predicted by Theorem 3.5.
2 Basic properties
We’ll make an extensive use of the first fundamental theorem of projective geometry (see [BergerMarcel1927-20162016GT1, Prop. 4.5.10]), which states that given two projective frames, there exists a unique projective transformation mapping the first to the second.
Lemma 2.1.
Polygonal transformations preserve monodromies.
Proof.
Let $P=((p_{i})_{i\in\mathbb{Z}},M)$ be an $n$-gon in $\mathbb{P}^{d}$, $T$ be a polygonal transformation, and $T(P)=:P^{\prime}=((p^{\prime}_{i})_{i\in\mathbb{Z}},M^{\prime})$. We have
$M\cdot P^{\prime}=((Mp^{\prime}_{i})_{i\in\mathbb{Z}},MM^{\prime}M^{-1}),$
and by the monodromy condition
$M\cdot P=((p_{i+n})_{i\in\mathbb{Z}},M).$
Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$
$M\cdot P^{\prime}=T(M\cdot P)=((p^{\prime}_{i+n})_{i\in\mathbb{Z}},M^{\prime}),$
so we get that for all $i\in\mathbb{Z}$
$M^{\prime}p^{\prime}_{i}=p^{\prime}_{i+n}=Mp^{\prime}_{i}.$
Since $M$ and $M^{\prime}$ map the projective frame $(p^{\prime}_{0},\dots,p^{\prime}_{d+1})$ to $(p^{\prime}_{n},\dots,p^{\prime}_{d+1+n})$, we conclude by the first fundamental theorem of projective geometry that they are equal. ∎
2.1 Coordinate system in the moduli space
First of all, we need a way to parameterize the moduli space. We do so by generalizing the usual cross-ratio in $\mathbb{P}^{1}$ to $\mathbb{P}^{d}$.
Definition 2.2.
Let $p_{0},\dots,p_{d+1}\in\mathbb{P}^{d}$ be points forming a projective frame. There exists a unique projective transformation $C_{p_{0},\dots,p_{d+1}}$ which maps $(p_{0},p_{1},\dots,p_{d+1})$ to the standard frame
$([1:\dots:1],[1:0:\dots:0],\dots,[0:\dots:0:1]).$
If $p$ is another point in $\mathbb{P}^{d}$, then its cross-ratio with respect to the frame $\bm{(p_{0},...,p_{d+1})}$ is
$c=[x_{1}:\dots:x_{d+1}]:=C_{p_{0},\dots,p_{d+1}}(p).$
The following proposition derives from the definition, and is proved exactly in the same way as for the $\mathbb{P}^{1}$ case (see [AudinMichèle2006G, Prop. 6.2] for details).
Proposition 2.3.
This cross-ratio is projectively invariant. Moreover, there exist a projective transformation mapping a frame $(p_{0},\dots,p_{d+1},p)$ to another one $(p_{0}^{\prime},\dots,p_{d+1}^{\prime},p^{\prime})$ if and only if their respective cross-ratios coincide.
There are explicit formulas involving determinants, using the so called volume cross-ratio defined in [Olver_2001]. For a geometric understanding involving only the usual cross-ratio on a projective line, see [BergerMarcel1927-20162016GT1, Prop. 6.5.11].
Proposition 2.4.
Let $p_{0},\dots,p_{d+1}$ be a projective frame in $\mathbb{P}^{d}$, $p\in\mathbb{P}^{d}$ be another point, and $V_{0},\dots,V_{d+1},V$ be a lift in $\mathbb{A}^{d+1}$. Let $x_{1}:\dots:x_{d+1}]$ be the cross ratio defined previously. For any $1\leq i<j\leq d+1$, call $V^{i,j}$ the collection of all vectors from $V_{1}$ to $V_{d+1}$, excluding $V_{i}$ and $V_{j}$. Then we have:
$\frac{x_{i}}{x_{j}}=\frac{\mathrm{det}(V^{i,j},V_{i},V_{0})}{\mathrm{det}(V^{i,j},V_{i},V)}\frac{\mathrm{det}(V^{i,j},V_{j},V)}{\mathrm{det}(V^{i,j},V_{j},V_{0})}.$
@@ 2.1 Coordinate system in the moduli space
This retrieves the usual definition of the cross-ratio of four points in $\mathbb{P}^{1}$:
$p_{1},p_{2},p_{0},p]=\frac{p_{1}-p_{0}}{p_{1}-p}\frac{p_{2}-p}{p_{2}-p_{0}}.$
Proof.
First of all, we see that any choice of lift gives the same quantity. Indeed, if any vector is multiplied by a non-zero scalar, it cancels out in the ratio of determinants. Then, we see it is projective invariant, as if we apply any linear automorphism $A$ to the vectors, all the determinants get multiplied by $\mathrm{det}(A)$, which in turn cancels out in the ratio. Hence, we can work with
$V_{0}=\begin{pmatrix}1\\ \vdots\\ 1\end{pmatrix},\quad V_{1}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix},\quad\dots,\quad V_{d+1}=\begin{pmatrix}0\\ \vdots\\ 0\\ 1\end{pmatrix},\quad V=\begin{pmatrix}x_{1}\\ \vdots\\ x_{d+1}\end{pmatrix}$.
We conclude by the straight out computation:
$\frac{\mathrm{det}(V^{i,j},V_{i},V_{0})}{\mathrm{det}(V^{i,j},V_{i},V)}\frac{\mathrm{det}(V^{i,j},V_{j},V)}{\mathrm{det}(V^{i,j},V_{j},V_{0})}=\frac{1\times x_{i}}{x_{j}\times 1}.\qed$